# Presto! Distilling Steps and Layers for Accelerating Music Generation

### v1

@@ Abstract
Despite advances in diffusion-based text-to-music (TTM) methods, efficient, high-quality generation remains a challenge. We introduce Presto!, an approach to inference acceleration for score-based diffusion transformers via reducing both sampling steps and cost per step. To reduce steps, we develop a new score-based distribution matching distillation (DMD) method for the EDM-family of diffus ion models, the first GAN-based distillation method for TTM. To reduce the cost per step, we develop a simple, but powerful improvement to a recent layer distillation method that improves learning via better preserving hidden state variance. Finally, we combine our step and layer distillation methods together for a dual-faceted approach. We evaluate our step and layer distillation methods independently and show each yield best-in-class performance. Our combined distillation method can generate high-quality outputs with improved diversity, accelerating our base model by 10-18x (230/435ms latency for 32 second mono/stereo 44.1kHz, 15x faster than comparable SOTA) — the fastest high-quality TTM to our knowledge.
@@ 1 Introduction
We have seen a renaissance of audio-domain generative media (Chen et al., 2024; Agostinelli et al., 2023; Liu et al., 2023; Copet et al., 2023), with increasing capabilities for both Text-to-Audio (TTA) and Text-to-Music (TTM) generation. This work has been driven in-part by audio-domain diffusion models (Song et al., 2020; Ho et al., 2020; Song et al., 2021), enabling considerably better audio modeling than generative adversarial network (GAN) or variational autoencoder (VAE) methods (Dhariwal & Nichol, 2021). Diffusion models, however, suffer from long inference times due to their iterative denoising process, requiring a substantial number of function evaluations (NFE) during inference (i.e. sampling) and resulting in $\approx$5-20 seconds at best for non-batched latency.
@@
Overall, our core contributions include the development of a holistic approach to accelerating score-based diffusion transformers including:
  • The development of distribution matching distillation for continuous-time score-based diffusion (i.e. EDM), the first GAN-based distillation method for TTA/TTM.
  • The development of an improved layer distillation method that consistently improves upon both past layer distillation method and our base diffusion model.
  • The development of the first combined layer and step distillation method.
  • Evaluation showing our step, layer, and layer-step distillation methods are all best-in-class and, when combined, can accelerate our base model by 10-18x (230/435ms latency for 32 second mono/stereo 44.1kHz, 15x faster than Stable Audio Open (Evans et al., 2024c)), the fastest TTM model to our knowledge. For sound examples, see https://presto-music.github.io/web/.
@@ 3.1 EDM-Style Distribution Matching Distillation
We first redefine DMD2 in the language of continuous-time, score-based diffusion models (i.e. EDM-style). Our goal is to distill our score model $\mu_{\bm{\theta}}$ (which we equivalently denote as $\mu_{\text{real}}$, as it is trained to model the score of real data) into an accelerated generator $G_{\bm{\phi}}$ that can sample in 1-4 steps. Formally, we wish to minimize the KL Divergence between the real distribution $p_{\text{real}}$ and the generator $G_{\bm{\phi}}$’s distribution $p_{\text{fake}}$: $\mathcal{L}_{\text{DMD}}=D(p_{\text{real}}\|p_{\text{fake}})$. The KL term cannot be calculated explicitly, but we can calculate its gradient with respect to the generator if we can access the score of the generator’s distribution. Thus, we also train a “fake” score model $\mu_{\bm{\psi}}$ (or equivalently, $\mu_{\text{fake}}$) to approximate the generator distribution’s score function at each gradient step during training.
@@ 3.2 Presto-S: Score-based Distribution Matching Distillation
Given the new notation, we proceed to develop our score-based distribution matching step distillation, Presto-S. Our algorithm is shown in Fig. 1, in Appendix 1, and discussed below.
@@ 3.3 Presto-L: Variance and Budget-Aware Layer Dropping
Given our step distillation approach above, we now seek methods to reduce the cost of individual steps themselves here through layer distillation, and then combine both step and layer distillation in Section 3.4. We begin with the current SOTA method: ASE (Moon et al., 2024). ASE employs a fixed dropping schedule that monotonically maps noise levels to compute budgets, allocating more layers to lower noise levels. We enhance this method in three key ways: (1) ensuring consistent variance in layer distilled outputs, (2) implementing explicit budget conditioning, and (3) aligning layer-dropped outputs through direct distillation.
@@ 4.1 Setup
Baselines: We compare against a number of acceleration algorithms and external models, including Consistency Models (CM) (Song et al., 2023), SoundCTM (Saito et al., 2024), DITTO-CTM (Novack et al., 2024a), DMD-GAN (Yin et al., 2024), ASE (Moon et al., 2024), MusicGen (Copet et al., 2023), and Stable Audio Open (Evans et al., 2024c). See Appendix A.2 for more details.
Metrics: We use Frechet Audio Distance (FAD) (Kilgour et al., 2018), Maximum Mean Discrepancy (MMD) (Jayasumana et al., 2024), and Contrastive Language-Audio Pretraining (CLAP) score (Wu et al., 2023), all with the CLAP-LAION music backbone (Wu et al., 2023) given its high correlation with human perception (Gui et al., 2024). FAD and MMD measure audio quality/realness with respect to Song Describer (lower better), and CLAP score measures prompt adherence (higher better). When comparing to other models, we also include density (measuring quality), recall and coverage (measuring diversity) (Naeem et al., 2020), and real-time factor (RTF) for both mono (M) and stereo (S, using MusicHiFi), which measures the total seconds of audio generated divided by the generation time, where higher is better for all.
@@ 4.3 Presto-L Results
We compare Presto-L with both our baseline diffusion model and ASE (Moon et al., 2024) using the 2nd order DPM++ sampler (Lu et al., 2022) with CFG++ (Chung et al., 2024). For ASE and Presto-L, we use the optimal “D3” configuration from Moon et al. (2024), which corresponds to a dropping schedule, in terms of decreasing noise level (in quintiles), of $[14,12,8,4,0]$ (i.e. we drop 14 layers for noise levels in the top quintile, 12 for the next highest quintile, and so on). Layer distillation results at various sampling budgets are shown in Fig. 6. Presto-L yields an improvement over the base model on all metrics, speeding up by $\approx$27and improving quality and text relevance. ASE provides similar acceleration but degrades performance at high sampling steps and scales inconsistently. The behavior of dropping layers improving performance can be viewed via the lens of multi-task learning, where (1) denoising each noise level is a different task (2) later layers only activating for lower noise levels enables specialization for higher frequencies. See Appendix A.7 for further ablations.
@@ A.1 Model Design Details
For Presto-S, following Yin et al. (2024) we use a fixed guidance scale of $w=4.5$ throughout distillation for the teacher model as CFG++ is not applicable for the distribution matching gradient. Additionally, we use a learning rate of 5e-7 with Adam. For all step distillation methods, we distill each model with a batch size of 80 across 16 Nvidia A100 GPUs for 32K iterations. We train all layer distillation methods for 60K iterations with a batch size of 12 across 16 A100 GPUs.
A.2 Baseline Description
We benchmark against multiple diffusion acceleration algorithms and open-source models:
@@ A.2 Baseline Description
  • SoundCTM (Saito et al., 2024): This approach distills a model into a consistency trajectory model (Kim et al., 2023) that enforces the self-consistency property for the entire diffusion process, learning an anywhere-to-anwhere mapping. SoundCTM forgoes the original CTM adversarial loss and calculates the consistency loss using intermediate features of the base model.
@@ A.3 Presto-S Algorithm
Our complete Presto-S algorithm is outlined below. For a visual representation, please see Fig. 1.
@@
A.4 Analyzing Failure Modes of Combined Layer and Step Distillation

### v2

@@ Abstract
Despite advances in diffusion-based text-to-music (TTM) methods, efficient, high-quality generation remains a challenge. We introduce Presto!, an approach to inference acceleration for score-based diffusion transformers via reducing both sampling steps and cost per step. To reduce steps, we develop a new score-based distribution matching distillation (DMD) method for the EDM-family of diffusion models, the first GAN-based distillation method for TTM. To reduce the cost per step, we develop a simple, but powerful improvement to a recent layer distillation method that improves learning via better preserving hidden state variance. Finally, we combine our step and layer distillation methods together for a dual-faceted approach. We evaluate our step and layer distillation methods independently and show each yield best-in-class performance. Our combined distillation method can generate high-quality outputs with improved diversity, accelerating our base model by 10-18x (230/435ms latency for 32 second mono/stereo 44.1kHz, 15x faster than the comparable SOTA model) — the fastest TTM to our knowledge.
@@ 1 Introduction
We have seen a renaissance of audio-domain generative media (Chen et al., 2024; Agostinelli et al., 2023; Liu et al., 2023; Copet et al., 2023), with increasing capabilities for both Text-to-Audio (TTA) and Text-to-Music (TTM) generation. This work has been driven in-part by audio-domain diffusion models (Song et al., 2020; Ho et al., 2020; Song et al., 2021), enabling considerably better audio modeling than generative adversarial network (GAN) or variational autoencoder (VAE) methods (Dhariwal & Nichol, 2021). Diffusion models, however, suffer from long inference times due to their iterative denoising process, requiring a substantial number of function evaluations (NFE) during inference (i.e. sampling) and resulting in $\approx$5-20 seconds at best for non-batched $\approx$32s outputs.
@@
Overall, our core contributions include the development of a holistic approach to accelerating score-based diffusion transformers including : (1) The development of distribution matching distillation for continuous-time score-based diffusion (i.e. EDM), the first GAN-based distillation method for TTM. (2) The development of an improved layer distillation method that consistently improves upon both past layer distillation method and our base diffusion model. (3) The development of the first combined layer and step distillation method. (4) Evaluation showing our step, layer, and layer-step distillation methods are all best-in-class and, when combined, can accelerate our base model by 10-18x (230/435ms latency for 32 second mono/stereo 44.1kHz, 15x faster than Stable Audio Open (Evans et al., 2024c)), the fastest TTM model to our knowledge. For sound examples (anonymous link), see https://presto-music.github.io/web/.
@@ 3.1 EDM-Style Distribution Matching Distillation
We first redefine DMD2 in the language of continuous-time, score-based diffusion models (i.e. EDM-style). Our goal is to distill our score model $\mu_{\bm{\theta}}$ (which we equivalently denote as $\mu_{\text{real}}$, as it is trained to model the score of real data) into an accelerated generator $G_{\bm{\phi}}$ that can sample in 1-4 steps. Formally, we wish to minimize the reverse KL Divergence between the real distribution $p_{\text{real}}$ and the generator $G_{\bm{\phi}}$’s distribution $p_{\text{fake}}$: $\mathcal{L}_{\text{DMD}}=D(p_{\text{fake}}\|p_{\text{real}})$. The KL term cannot be calculated explicitly, but we can calculate its gradient with respect to the generator if we can access the score of the generator’s distribution. Thus, we also train a “fake” score model $\mu_{\bm{\psi}}$ (or equivalently, $\mu_{\text{fake}}$) to approximate the generator distribution’s score function at each gradient step during training.
@@ 3.2 Presto-S: Score-based Distribution Matching Distillation
We develop our score-based distribution matching step distillation, Presto-S below and in Fig. 1 as well as the algorithm in Appendix A.3, a pseudo-code walkthrough in Appendix A.4, and expanded visualization in Appendix A.5.
@@ 3.3 Presto-L: Variance and Budget-Aware Layer Dropping
Given our step distillation approach above, we now seek to reduce the cost of individual steps themselves through layer distillation, and then combine both step and layer distillation in Section 3.4. We begin with the current SOTA method: ASE (Moon et al., 2024). ASE employs a fixed dropping schedule that monotonically maps noise levels to compute budgets, allocating more layers to lower noise levels. We enhance this method in three key ways: (1) ensuring consistent variance in layer distilled outputs, (2) implementing explicit budget conditioning, and (3) aligning layer-dropped outputs through direct distillation. See Appendix A.6 for more details.
@@ 4.1 Setup
Baselines: We compare against a number of acceleration algorithms using our base model: Consistency Models (CM) (Song et al., 2023), SoundCTM (Saito et al., 2024), DITTO-CTM (Novack et al., 2024a), DMD-GAN (Yin et al., 2024), and ASE (Moon et al., 2024), as well as MusicGen (Copet et al., 2023) and Stable Audio Open (Evans et al., 2024c). See Appendix A.2 for more details.
Metrics: We use a number of common evaluation metrics for text-to-music generation, including distributional quality/diversity metrics (FAD/MMD/Density/Recall/Coverage), prompt adherence (CLAP Score), and latency (RTF). See Appendix A.2 for more details.
@@ 4.3 Presto-L Results
We compare Presto-L with both our baseline diffusion model and ASE (Moon et al., 2024) using the 2nd order DPM++ sampler (Lu et al., 2022) with CFG++ (Chung et al., 2024). For ASE and Presto-L, we use the optimal “D3” configuration from Moon et al. (2024), which corresponds to a dropping schedule, in terms of decreasing noise level (in quintiles), of $[14,12,8,4,0]$ (i.e. we drop 14 layers for noise levels in the top quintile, 12 for the next highest quintile, and so on). Layer distillation results at various sampling budgets are shown in Fig. 6. Presto-L yields an improvement over the base model on all metrics, speeding up by $\approx$27and improving quality and text relevance. ASE provides similar acceleration but degrades performance at high sampling steps and scales inconsistently. Dropping layers improving performance can be viewed via the lens of multi-task learning, where (1) denoising each noise level is a different task (2) later layers only activating for lower noise levels enables specialization for higher frequencies. See Appendix A.10 for further ablations.
@@ Acknowledgements
Ethics Statement and Reproducibility
As TTM systems become more powerful, there is both the opportunity to increase accessibility of musical expression, but also concern such systems may compete with creators. To reduce risk, we train our TTM work only on instrumental licensed music. Additionally, we hope that our focus on efficiency is useful to eventually make interactive-rate co-creation tools, allowing for greater flexibility and faster ideation. Following these concerns, we do not plan to release our model, but have done our best to compare against multiple open source baselines and/or re-train alternative methods for comparison and in-depth understanding of the reproducible insights of our work.
@@ A.1 Model Design Details
Our discriminator design follows Yin et al. (2024) with a number of small modifications. $D_{\bm{\psi}}$ consists of 4 blocks of 1D convolutions interleaved with GroupNorm and SiLU activations, and a final linear layer to collapse the channel dimension. The discriminator thus does not use any final linear layer to project to a single value, and instead its’ output is also a 1D sequence but at even heavier downsampling than the input representation at $\approx$2.8 Hz. The discriminator receives its’ input from the output of the 14th DiT Block (i.e. the halfway point through our 28 block DiT), as DiTs lack a clear “bottleneck” layer to place the discriminator like in UNets. We leave further investigation into discriminator design and placement inside the model for future work.
@@
For Presto-S, following Yin et al. (2024) we use a fixed guidance scale of $w=4.5$ throughout distillation for the teacher model as CFG++ is not applicable for the distribution matching gradient. We use 5 fake score model (and discriminator) updates per generator update, following Yin et al. (2024), as we found little change in performance when varying the quantity around 5 (though using $\leq 3$ updates resulted in large training instability). Note that throughout Presto-S, the fake score model and the discriminator share an optimizer state. Additionally, we use a learning rate of 5e-7 with Adam for both the generator and fake score model / discriminator. We set $\nu_{1}=0.01$ and $\nu_{2}=0.005$ following Yin et al. (2024). For all step distillation methods, we distill each model with a batch size of 80 across 16 Nvidia A100 GPUs for 32K iterations. We train all layer distillation methods for 60K iterations with a batch size of 12 across 16 A100 GPUs with a learning rate of 8e-5. For Presto-L, we set $\nu=0.1$.
A.2 Experimental Details
A.2.1 Baseline Details
Our benchmarks are divided into two main classes: acceleration algorithms and external open-source models. For acceleration algorithms, we distill our internal base model per method, utilizing publicly available code as a reference when available (Song et al., 2023; Saito et al., 2024; Yin et al., 2024). For the open-source external models, we use the models directly in their default setups as recommended by Copet et al. (2023); Evans et al. (2024c).
@@ A.2.1 Baseline Details
  • SoundCTM (Saito et al., 2024): This approach distills a model into a consistency trajectory model (Kim et al., 2023) that enforces the self-consistency property, learning an anywhere-to-anywhere mapping. SoundCTM forgoes the original CTM adversarial loss and calculates the consistency loss via intermediate base model features.
@@
A.2.2 Metrics Details
We use Frechet Audio Distance (FAD) (Kilgour et al., 2018), Maximum Mean Discrepancy (MMD) (Jayasumana et al., 2024), and Contrastive Language-Audio Pretraining (CLAP) score (Wu et al., 2023), all with the CLAP-LAION music backbone (Wu et al., 2023) given its high correlation with human perception (Gui et al., 2024). FAD and MMD measure audio quality/realness with respect to Song Describer (lower better), and CLAP score measures prompt adherence (higher better). When comparing to other models, we also include density (measuring quality), recall and coverage (measuring diversity) (Naeem et al., 2020), and real-time factor (RTF) for both mono (M) and stereo (S, using MusicHiFi), which measures the total seconds of audio generated divided by the generation time, where higher is better for all.
@@ A.3 Presto-S Algorithm
We outline a condensed algorithm of Presto-S in math notation in Algorithm 1.
A.4 Presto-S Pseudo-code Walkthrough
We provide a comprehensive algorithm walkthrough using PyTorch psuedo-code of our Presto-S training loop below. To perform Presto-S, we first define the corruption process for any given clean sample, according to either the training $p(\sigma^{\text{train}})$ or the inference $p(\sigma^{\text{inf}})$ noise distribution:
1def diffuse(x, dist):
2 eps = noise_normal_like(x)
3 if dist == ’training’:
4 sigma = training_dist_like(x)
5 elif dist == ’inference’:
6 sigma = inference_dist_like(x)
7 return x + sigma * eps, sigma
We then define each of the component loss functions for the Presto-S continuous-time DMD2 distillation process. This corresponds to the three loss types: the distribution matching loss, the least-squares GAN loss, and the fake denoising score matching loss. For the distribution matching loss, we corrupt some generated sample according to the training distribution and then pass that into both the fake and real score models (where the real score model uses classifier-free guidance). The difference in these scores forms the distribution matching gradient:
1def dmd(x, real_score_model, fake_score_model, cfg):
2 x_noise, sigma = diffuse(x, ’training’)
3 fake_denoised = fake_score_model(x_noise, sigma)
4 real_denoised = real_score_model(x_noise, sigma, cfg)
5 return fake_denoised - real_denoised
For the least-squares GAN loss, we corrupt some sample (either real or generated) according to the training distribution and pass this through the discriminator (which itself involves first passing through some of the fake score model to extract intermediate features). The output of the discriminator is then passed into the least-squares loss against some target value (i.e. the generator wants to push the discriminator outputs on generated samples towards 1, while the discriminator aims to push generated samples towards 0 and real samples towards 1):
1def gan(x, discriminator, tgt=1):
2 x_noise, sigma = diffuse(x, ’training’)
3 d_out = discriminator(x_noise, sigma)
4 return mse(tgt, d_out)
Finally, we have the fake DSM loss. This loss is identical to the normal diffusion loss (with a weighted MSE between the outputs of the score model and the clean data), yet will be calculated treating generator outputs as the ground truth clean data and using the fake score model:
1def dsm(x, fake_score_model):
2 x_noise, sigma = diffuse(x, ’training’)
3 x_denoised = fake_score_model(x_noise, sigma)
4 return weighted_mse(x, x_denoised, sigma)
Given these helper loss functions, we can now proceed with the main distillation loop, which is as follows. For both the generator and discriminator turns, we first corrupt some real input data according to the inference distribution, and pass this through our generator to get the generator outputs x_denoised (steps (1) and (4) in Fig. 8). If it is a generator turn (which happens once for every 5 fake score turns), we calculate the distribution matching loss (step (2)) and the generator adversarial loss (step (3)) on x_denoised and update the generator. If it is a fake score turn, we calculate and the discriminator’s adversarial loss (step (5)) on both the generated x_denoised and real samples x and the fake DSM loss (step (6)) on x_denoised, thus updating the fake score model and the discriminator:
1def forward(
2 x, generator, discriminator, fake_score_model, real_score_model, generator_turn, nu_1, nu_2
3):
4 # step (1) and (4)
5 x_noise, sigma = diffuse(x, ’inference’)
6 x_denoised = generator(x_noise, sigma)
7
8 if generator_turn: # GENERATOR TURN
9 # Distribution Matching Loss, step (2)
10 dmd_loss = dmd(x_denoised, real_score_model, fake_score_model, cfg)
11
12 # Generator Adversarial Loss, step (3)
13 g_loss = gan(x_denoised, discriminator, 1)
14
15 loss = dmd_loss + nu_1 * g_loss
16 else: # FAKE SCORE TURN
17 # Discriminator Adversarial Loss, step (5)
18 d_loss = gan(x, discriminator, 1) + gan(x_denoised, discriminator, 0)
19
20 # fake DSM loss, step (6)
21 dsm_loss = dsm(x_denoised, fake_score_model)
22
23 loss = dsm_loss + nu_2 * d_loss
24 return loss
This constitutes one full update of the Presto-S process, alternating between the generator and fake score model / discriminator updates. At inference time, we can feed in pure noise and alternate between generating clean data with our generator and adding progressively smaller noise back to the generation (for some pre-defined list of noise levels), allowing for multi-step sampling:
1def inference(generator, sigmas, start_noise):
2 x = start_noise
3 for sigma in sigmas:
4 x = x + noise_normal_like(x) * sigma
5 x = generator(x, sigma)
6 return x
A.5 Presto-S Expanded Diagram
For an in-depth visual illustration of Presto-S, please see  Fig. 8 and Fig. 9 for expanded training and inference diagrams.
Figure 8: Presto-S training process.
Figure 9: Presto-S inference. For multi-step sampling, we use ping-pong-like sampling.
A.6 Presto-L Algorithm
Algorithm 2 Presto-L
0:  : pre-trained score model $\mu_{\bm{\theta}}$, real sample $\bm{x}_{\text{real}}$, self-teacher weight $\nu$, optimizer $g,g_{2}$, weighting function $\lambda$, # of DiT blocks $B$, budget mapping $\ell$, layer drop function $\mathbf{LD}$
1:  $\sigma\sim p(\sigma^{\text{train}})$
2:  $b=\ell(\sigma)$
3:  $\epsilon\sim\mathcal{N}(0,\bm{I})$
4:  $\hat{\bm{x}}_{\bm{L}},h_{\bm{L}}=\mathbf{LD}(\mu_{\bm{\theta}},\bm{x}_{\text{real}}+\sigma\epsilon,\sigma,b)$
5:  $\hat{\bm{x}}_{\text{full}},h_{\text{full}}=\mathbf{LD}(\mu_{\bm{\theta}},\bm{x}_{\text{real}}+\sigma\epsilon,\sigma,B)$
6:  $\mathcal{L}_{\text{DSM}}=\lambda(\sigma)\|\bm{x}_{\text{real}}-\hat{\bm{x}}_{\bm{L}}\|_{2}^{2}$
7:  $\mathcal{L}_{\text{st}}=\|h_{\bm{L}}-\texttt{sg}(h_{\text{full}})\|_{2}^{2}$
8:  $\bm{\theta}\leftarrow\bm{\theta}-g(\nabla_{\bm{\theta}}\mathcal{L}_{\text{DSM}}+\nu\nabla_{\bm{\theta}}\mathcal{L}_{\text{st}})$
8:  : $\bm{\theta}$
Algorithm 3 $\mathbf{LD}$: Modified DiT forward pass with layer dropping and budget conditioning.
0:  : score model noise embedder $\mu_{\bm{\theta}}^{\text{noise}}$, score model budget embedder $\mu_{\bm{\theta}}^{\text{budget}}$, score model DiT blocks $\{\mu_{\bm{\theta}}^{i}\}_{i=1}^{B}$, score model budget AdaLN $\mu_{\bm{\theta}}^{\text{LN}}$, score model output layer $\mu_{\bm{\theta}}^{\text{final}}$, input $\bm{x}$, noise level $\sigma$, budget $b$
1:  $\bm{e}_{\sigma}=\mu_{\bm{\theta}}^{\text{noise}}(\sigma)$ // embed noise level
2:  $\bm{e}_{b}=\mu_{\bm{\theta}}^{\text{budget}}(b)$ // embed budget
3:  $\bm{e}=\bm{e}_{\sigma}+\bm{e}_{b}$
4:  for  $i:=1$ to $b-1$ do
5:     // apply first b-1 DiT blocks
6:     $x=\mu_{\bm{\theta}}^{i}(x,\bm{e})$
7:  end for
8:  $x=\mu_{\bm{\theta}}^{B}(x,\bm{e})$ // apply final DiT block
9:  $x=\mu_{\bm{\theta}}^{\text{LN}}(x,\bm{e}_{b})$ // apply budget-based AdaLN
10:  $h=x/\|x\|_{2}$ // get normalized hidden state for $\mathcal{L}_{\text{st}}$
10:  : $\mu_{\bm{\theta}}^{\text{final}}(x),h$
We show the full algorithm in detail for Presto-L in Algorithm 2, which proceeds as a modified version of standard diffusion training like in Moon et al. (2024). We first sample some noise level $\sigma$, and then map the noise level to its corresponding budget $b$ given some mapping function $\ell(\cdot)$. Following Moon et al. (2024), :$\ell:\mathbb{R}\rightarrow\{i\}_{i=1}^{B}$ is a deterministic map from the percentile of the noise level according to the training noise distribution $F(\sigma)$ (where $F$ is the cumulative distribution function) to some budget amount, which we write as $[q_{1},q_{2},q_{3},q_{4},q_{5}]$ for a mapping based on descending quintiles (e.g. $q_{1}=14$ means that all noise levels in the largest quintile drop 14 layers).
We then call the modified forward function of the model $\mathbf{LD}$ (see Algorithm 3) on the noisy inputs with both the given budget $b$ and the full budget $B$ (i.e. using all DiT blocks). $\mathbf{LD}$ modifies the forward pass of the model by (1) adding a global budget embedding that is added to the noise embedding (2) only iterating through the first $b-1$ DiT blocks followed by the final DiT block (to preserve final block behavior, see Section 3.3) (3) adding an additional AdaLN conditional only on the budget after the final DiT block and (4) also returning the normalized hidden state of the model (i.e. the input to the final layer of the DiT, normalized along the channel dimension). We calculate the standard denoising score matching loss $\mathcal{L}_{\text{DSM}}$ as normal but with our layer-dropped outputs, and additionally calculate $\mathcal{L}_{\text{st}}$ as the MSE between the layer-dropped hidden state and the full budget hidden state (with a stop-gradient operation on the full budget pass.
A.7 Analyzing Failure Modes of Combined Layer and Step Distillation